An object is accelerating at a. Select all of the following actions that would cause the object to accelerate at 2 Check all that apply. Halve the net force while keeping mass constant. Double the net force while keeping mass constant. Double the net force and quadruple the mass. Double the mass while keeping net force constant. Halve the mass while keeping net force constant..
Answer #1

true, since force is directly proportional to Acceleration.

======

false.

=======

true,

a2 / a1 = m1 F2 / ( m2 F1 )

a2 / a = m1 ( 2 F1) / ( 4 m1 F1)

a2 = a/2

========

true, mass is inversely proportional to Acceleration

======

false

=======

Comment in case any doubt, will reply for sure.. Goodluck